OpenStudy (anonymous):

\[\text{find }n\text{ for wich the expression}\\2^n+3^n \text{ is divisible by} 13\]

OpenStudy (anonymous):

\[\large 13|2^n+3^n\\ 13|2^n ,13|3^n\]

OpenStudy (blockcolder):

That's not necessarily true. For instance, 2|(1+3) but 2 doesn't divide either 1 or 3.

OpenStudy (anonymous):

so we need to have \[\large 2^n\equiv 0\mod13\]

OpenStudy (anonymous):

look at the question

OpenStudy (anonymous):

we r looking for n that will make this true

OpenStudy (anonymous):

@mukushla

ganeshie8 (ganeshie8):

2 wil make it true, but 13 wont divide 2^2 or 3^2

ganeshie8 (ganeshie8):

i think your start is a bad one

OpenStudy (anonymous):

but we r looking for a scenario where it divides them

OpenStudy (anonymous):

okay number theory???

OpenStudy (anonymous):

approach

ganeshie8 (ganeshie8):

n = 2 will make the equation 2^n + 3^n = 13k true right ? i want u pause and check ur start

ganeshie8 (ganeshie8):

oh @mukushla is here :)

OpenStudy (blockcolder):

I found that n=2, 6, and 10 work, so a conjecture could be made that n=4k+2 where k is an integer will work. Also, this is provable via induction on k. Inductive step: Assume that \[2^{4k+2}+3^{4k+2}\equiv0\pmod{13}\] We must prove that \[2^{4(k+1)+2}+3^{4(k+1)+2}\equiv0\pmod{13}\] To prove this, we have \[\begin{align} 2^{4(k+1)+2}+3^{4(k+1)+2}&\equiv2^{4k+4+2}+3^{4k+4+2}\pmod{13}\\ &\equiv2^{4k+2}+3^{4k+2}+16+81\pmod{13}\\ &\equiv 2^{4k+2}+3^{4k+2}+97\pmod{13}\\ &\equiv 2^{4k+2}+3^{4k+2}\pmod{13}\text{ because }13|97\\ &\equiv 0\pmod{13}\text{ by inductive assumption} \end{align}\] Thus, if n=4k+2, where k is an integer, then 2^n+3^n will divide 13.

OpenStudy (anonymous):

okay i see ,but is the any other numbers ...say less than hundred,\[1\len\le 100\]

OpenStudy (anonymous):

\[1\le n \le100\]

OpenStudy (blockcolder):

Just plug in k=1, 2, ..., 24 (k=24 gives n=98) and you get all n where 13 divides 2^n+3^n, and n is between 1 and 100.

OpenStudy (anonymous):

so a follow up question says.hence prove that \[13|2^{50}+3^{50}\]

OpenStudy (anonymous):

thanks at @blockcolder for the first part

OpenStudy (blockcolder):

50=4(12)+2, so simply let k=12.

OpenStudy (anonymous):

lets say we didnt do the induction how can we do this from scratch

OpenStudy (anonymous):

wat steps are necesarry to show this

OpenStudy (anonymous):

by fermat \[2^{12}\equiv 1\mod 13\]

OpenStudy (anonymous):

\[\large 2^{50}=2^{2(12)}2^2\]